OpenStudy (anonymous):

wait like when you have (y^5)^9 you just multiply 5 AND 9?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yes but you must square them

OpenStudy (anonymous):

= y ^(5*9) = y^45

OpenStudy (anonymous):

kk thanks much :)